a child's money box contains 100 coins some of which are 5 cents and the rest 10 cent coins with a total value of ksh 6 .how many coins of each type does the box contain?
Answer by Boreal(15235)

5 cents=x
10 cents=100-x
x*.05+(100-x)*.10=6.00
.05x+10-.10x=6.00
-.05x=-4.00
x=80 5 cent
100-x=20 10 cent